OpenStudy (anonymous):

solve the trinomial r2 − 10r + 24 = 0

OpenStudy (unklerhaukus):

\[ax^2+bx+c=0\] \[x_{1,2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (unklerhaukus):

can you see what \(a,b,c\) are?

hartnn (hartnn):

or u can factorize it. what are the 2 numbers whose product is 24 and sum is -10. hint: both numbers are negative.